# DriveMind: A Dual-VLM based Reinforcement Learning Framework for Autonomous Driving

### v1

@@ C.4 Implementation Details
All components of DriveMind are implemented in PyTorch and seamlessly integrated with Stable-Baselines3 to leverage its robust Soft Actor-Critic (SAC) implementation, efficient replay buffers, and distributed training utilities. Our project is organized as a modular codebase with clearly separated packages for perception, semantic reasoning, world modeling, reinforcement learning, and utilities. The perception module contains the BEV projection pipeline: raw camera images and LiDAR point clouds are first preprocessed using a custom C++/CUDA extension to accelerate the top-down projection, after which a lightweight semantic segmentation network (based on DeepLabV3) produces multi-channel BEV masks for lanes, vehicles, pedestrians, and static obstacles. These BEV masks are normalized and resized to $224\times 224$ pixels and passed through the frozen CLIP ViT-bigG-14 image encoder to produce the static embedding $v_{t}$. The semantic_reasoning package wraps both the frozen contrastive static VLM and the dynamic encoder-decoder VLM (SmolVLM); it manages the novelty buffer $\mathcal{B}$, computes the $L_{2}$ novelty score $\Delta_{t}$, and handles asynchronous prompt generation via a background thread to avoid blocking the main control loop. Chain-of-thought distillation logic resides in a dedicated trainer class that performs an online per-sample distillation using output generated from GPT-4o-mini API, also handling text tokenization, teacher, student loss computation, and periodic checkpointing. The world_model package implements the compact MLP predictor $M_{\phi}$, complete with custom loss wrappers to compute prediction error and integrate it with the RL replay buffer for joint sampling. Finally, the rl_agent module configures the SAC actor and critic networks, each comprising a shared feature extractor that fuses BEV embeddings with ego-state and waypoint embeddings, alongside target network management, entropy tuning, and optimizer hooks. All modules share a common configuration system based on YAML files, allowing reproducible experiments via versioned config snapshots. Logging and visualization employ TensorBoard and Weights & Biases, capturing per-step metrics, network gradients, and video rollouts for qualitative inspection.
@@
Actor update frequency $2$ steps
Critic target update frequency $2$ steps
@@
Chain-of-Thought Distillation and Dynamic VLM Fine-Tuning.
The dynamic VLM, SmolVLM-256M, is fine-tuned to generate “present” and “ideal” prompts along with chain-of-thought (CoT) rationales based on novelty-triggered events. For each BEV image tensor captured during driving, a GPT-4o-mini API as a teacher is queried with a fixed structured prompt to generate CoT outputs, including a risk assessment, a current scene description, and future guidance, serialized in JSON format. The agent processes single samples on-the-fly: each incoming image is paired with the teacher’s labels to create a chat-formatted input that is used for lightweight adaptation of SmolVLM using Low-Rank Adaptation (LoRA).
@@
As summarized in Table 5, all experiments were conducted on a dual-socket AMD EPYC 7413 server, with each socket providing 24 physical cores (48 hardware threads) clocked up to 2.65 GHz and backed by 128 GB of DDR4-3200 memory in a quad-channel configuration. The system’s PCIe 4.0 lanes deliver high-bandwidth connectivity to a pair of NVIDIA RTX 3070 GPUs, each equipped with 24 GB of GDDR6 memory. One GPU was dedicated to hosting the CARLA simulation server (including sensor rendering and traffic management), while the second handled all reinforcement learning workloads, world-model ensemble training, and on-demand VLM prompt generation. The host runs Ubuntu 20.04.6 LTS (kernel 5.4.0-172), NVIDIA driver 470.239.06, and CUDA 11.4, with our deep learning stack built on Python 3.10, PyTorch 2.0 (using the CUDA 11.7 runtime), and NCCL for multi-GPU communication. Reinforcement learning algorithms (SAC and PPO) are implemented via the Stable-Baselines3 library, with mixed-precision support provided by Apex to accelerate training. Semantic reasoning leverages a frozen CLIP ViT-bigG-14 model alongside a 256 M-parameter SmolVLM fine-tuned via chain-of-thought distillation using GPT-4, both served from host RAM to minimize I/O overhead. Under these conditions, a full $10^{6}$-step DriveMind training run (including policy updates, world-model optimization, and novelty-triggered VLM decoding) requires approximately 56 wall-clock hours when distributed across the two GPUs. All baseline methods were evaluated on the identical hardware and software configuration to ensure a fair and reproducible comparison.
@@ C.5 Training Results
Table 6: Per-step latency breakdown on dual RTX 3070 (in milliseconds, two decimal places).
@@
Prompt-cache lookup 0.11
World-model predictive foresight 0.53
Baseline per-step total (no dynamic VLM) 9.06
@@
Overall per-step latency (amortized) 38.81

### v2

@@ C.4 Implementation Details
All components of DriveMind are implemented in PyTorch and seamlessly integrated with Stable‐Baselines3 to leverage its robust Soft Actor-Critic (SAC) implementation, efficient replay buffers, and distributed training utilities. Our project is organized as a modular codebase with clearly separated packages for perception, semantic reasoning, world modeling, reinforcement learning, and utilities. The perception module contains the BEV projection pipeline: raw camera images and LiDAR point clouds are first preprocessed using a custom C++/CUDA extension to accelerate the top‐down projection, after which a lightweight semantic segmentation network (based on DeepLabV3) produces multi‐channel BEV masks for lanes, vehicles, pedestrians, and static obstacles. These BEV masks are normalized and resized to $224\times 224$ pixels and passed through the frozen CLIP ViT‐bigG‐14 image encoder to produce the static embedding $v_{t}$. The semantic_reasoning package wraps both the frozen contrastive static VLM and the dynamic encoder-decoder VLM (SmolVLM); it manages the novelty buffer $\mathcal{B}$, computes the $L_{2}$ novelty score $\Delta_{t}$, and handles asynchronous prompt generation via a background thread to avoid blocking the main control loop. Chain‐of‐thought distillation logic resides in a dedicated trainer class that performs an online per-sample distillation using output generated from GPT‐4o-mini API, also handling text tokenization, teacher, student loss computation, and periodic checkpointing. The world_model package implements the compact MLP predictor $M_{\phi}$, complete with custom loss wrappers to compute prediction error and integrate it with the RL replay buffer for joint sampling. Finally, the rl_agent module configures the SAC actor and critic networks, each comprising a shared feature extractor that fuses BEV embeddings with ego‐state and waypoint embeddings, alongside target network management, entropy tuning, and optimizer hooks. All modules share a common configuration system based on YAML files, allowing reproducible experiments via versioned config snapshots. Logging and visualization employ TensorBoard and Weights & Biases, capturing per‐step metrics, network gradients, and video rollouts for qualitative inspection.
@@
Actor update frequency 22 steps
Critic target update frequency 22 steps
@@
Chain‐of‐Thought Distillation and Dynamic VLM Fine‐Tuning.
The dynamic VLM, SmolVLM‐256M, is fine‐tuned to generate “present” and “ideal” prompts along with chain‐of‐thought (CoT) rationales based on novelty-triggered events. For each BEV image tensor captured during driving, a GPT-4o-mini API as a teacher is queried with a fixed structured prompt to generate CoT outputs, including a risk assessment, a current scene description, and future guidance, serialized in JSON format. The agent processes single samples on-the-fly: each incoming image is paired with the teacher’s labels to create a chat-formatted input that is used for lightweight adaptation of SmolVLM using Low-Rank Adaptation (LoRA).
@@
As summarized in Table 5, all experiments were conducted on a dual‐socket AMD EPYC 7413 server, with each socket providing 24 physical cores (48 hardware threads) clocked up to 2.65 GHz and backed by 128 GB of DDR4‐3200 memory in a quad‐channel configuration. The system’s PCIe 4.0 lanes deliver high‐bandwidth connectivity to a pair of NVIDIA RTX 3070 GPUs, each equipped with 24 GB of GDDR6 memory. One GPU was dedicated to hosting the CARLA simulation server (including sensor rendering and traffic management), while the second handled all reinforcement learning workloads, world‐model ensemble training, and on-demand VLM prompt generation. The host runs Ubuntu 20.04.6 LTS (kernel 5.4.0-172), NVIDIA driver 470.239.06, and CUDA 11.4, with our deep learning stack built on Python 3.10, PyTorch 2.0 (using the CUDA 11.7 runtime), and NCCL for multi-GPU communication. Reinforcement learning algorithms (SAC and PPO) are implemented via the Stable-Baselines3 library, with mixed‐precision support provided by Apex to accelerate training. Semantic reasoning leverages a frozen CLIP ViT-bigG-14 model alongside a 256 M-parameter SmolVLM fine-tuned via chain-of-thought distillation using GPT-4, both served from host RAM to minimize I/O overhead. Under these conditions, a full $10^{6}$-step DriveMind training run (including policy updates, world-model optimization, and novelty-triggered VLM decoding) requires approximately 56 wall-clock hours when distributed across the two GPUs. All baseline methods were evaluated on the identical hardware and software configuration to ensure a fair and reproducible comparison.
@@ C.5 Training Results
Table 6: Per‐step latency breakdown on dual RTX 3070 (in milliseconds, two decimal places).
@@
Prompt‐cache lookup 0.11
World‐model predictive foresight 0.53
Baseline per‐step total (no dynamic VLM) 9.06
@@
Overall per‐step latency (amortized) 38.81